OpenStudy (anonymous):

Solve x^2 + 9x + 20 = 0. x = 2, x = 10 x = -2, x = -10 x = -5, x = -4 x = 5, x = 4

OpenStudy (anonymous):

@bibby @Haseeb96

OpenStudy (haseeb96):

X^2+4x+5x+20=0 x(x+4)5(x+4) (x+4)(x+5)=0 Either x+4=0. Or. x+5=0 x=-4. Or. x=-5

OpenStudy (anonymous):

So C?

OpenStudy (haseeb96):

Yes.

OpenStudy (anonymous):

Thanks